OpenStudy (anonymous):

Integrate h(x, y) = yi + xj over the circle of radius 1 centered at the origin traversed counterclockwise. a)1 b)0 c)pi d)-1 e)2pi f)N/A

OpenStudy (anonymous):

I think so, but I'm not sure/ This deals with line integrals

OpenStudy (anonymous):

thats actually from the problem. i don't really know what to do about the circle with radius 1

OpenStudy (anonymous):

the answer is 0, according to my textbook. I don't know why though

OpenStudy (anonymous):

You're asked to find \[\int_C {\bf h}(x,y)\cdot{\bf r}~d{\bf r}\] where \(C\) is the path of the circle, \({\bf h}(x,y)\) is the given vector field, \(y~\vec{i}+x~\vec{j}\), and \({\bf r}\) is the parameterization for the curve \(C\). You can go ahead with setting up the line integral, or you can check to see if the vector field is conservative. Conservative vector fields make any line integral over any closed path in the field zero.

OpenStudy (anonymous):

so @SithsAndGiggles how do you parametrize r? Because a circle is x^2+y^2=r^2, would the parameterization be u^2+u^2 ?

OpenStudy (anonymous):

Not quite, you'll want to use a different parameter for \(x\) and \(y\). Given a circle \(x^2+y^2=r^2\), you can express it in terms of the Pythagorean identity by setting \(x=r\cos t\) and \(y=r\sin t\), since \[x^2+y^2=r^2\cos^2t+r^2\sin^2t=r^2\] with \(t\) ranging from \(0\) to \(2\pi\). I recommend checking to see if the vector field is conservative first, though setting up the line integral is still good practice.

OpenStudy (anonymous):

I'm pretty sure the "vector field" is conservative. So r(u)=\[r^{2}\cos^2 (u)+r ^{2}\sin^2(u)?\]

OpenStudy (anonymous):

i'm using u because that's how our book does it; with r(u)

OpenStudy (anonymous):

You want \({\bf r}(u)\) to be a vector-valued function, so you should have \[{\bf r}(u)=\langle \cos u,~\sin u\rangle=\cos u~\vec{i}+\sin u~\vec{j}\\ d{\bf r}(u)=-\sin u~du~\vec{i}+\cos u~du~\vec{j}\] and using this parameterization means \[{\bf h}(x(u),y(u))=\langle \cos u,~\sin u\rangle=\cos u~\vec{i}+\sin u~\vec{j}\] as well. Since \(0\le u\le2\pi\), you have \[\int_C{\bf h}\cdot{\bf r}~d{\bf r}=\int_0^{2\pi}(\cos u~\vec{i}+\sin u~\vec{j})\cdot(-\sin u~\vec{i}+\cos u~\vec{j})~du\]

OpenStudy (anonymous):

wow that looks hard.

OpenStudy (anonymous):

ok so following your method, i got 0.

OpenStudy (anonymous):

alright thanks @SithsAndGiggles

OpenStudy (anonymous):

You're welcome. The easier way would be to check the partial derivatives of the vector field components. If \({\bf h}(x,y)=M(x,y)\vec{i}+N(x,y)\vec{j}\), then the vector field is conservative if \[\frac{\partial M}{\partial y}=\frac{\partial N}{\partial x}\] which is true in this case, since both partials come out to zero.

OpenStudy (anonymous):

alright @SithsAndGiggles . good to know

OpenStudy (zarkon):

The only thing I would add is to be mindful of your notation \[\int_C \mathbf{h}(\mathbf{r})\cdot\,d\mathbf{r} = \int_a^b \mathbf{h}(\mathbf{r}(u))\cdot\mathbf{r}'(u)\,du.\]

OpenStudy (anonymous):

thanks @Zarkon. I have that formula in my notes too.